There is a close "Big Ben" in London Fasting 60.0 kg, the long needle is 4.50 m long and weighs 100.0 kg. It has a rotating axis The kinetic love of love can go into circulation for an impossible two years. The left face is scaly Consider a long narrow rod with a distributed mass.
1
Expert's answer
2020-05-15T16:25:24-0400

Solution.

"m_1=60kg;"

"m_2=100kg;"

"l=4.50m;"

"\upsilon=\dfrac{2\pi R}{T};"

The period of rotation of the needle:

"T=12 hr=43200 s;"

Needle rotation speed:

"v=\dfrac{2\sdot3.14\sdot4.50m}{43200s}=0.0006541m/s;"


Kinetic energy of needle movement:

"W_k=\dfrac{mv^2}{2};"

"W_k=\dfrac {160kg\sdot( 0.0006541m/s)^2}{2}=0.0000342J;"

Answer: "W_k=0.00003442J."